1.1 x 1010) SOA ao points) Calculate the molar solubility of Baso a CK for Ba a) (5 points) in water b) (5 points in a solution containing 2.00 M so 2
Answer #1

BaSO4 <--> Ba+2 + SO4-2

Ksp = [BA+2][SO4-2]

a)

in water

Ksp = [BA+2][SO4-2]

1.1*10^-10 = S*S

S = sqrt(1.1*10^-10) = 1.05*10^-5 M

b)

in [SO4-2] = 2 M

so

Ksp = [BA+2][SO4-2]

1.1*10^-10 = [Ba+2](2)

[Ba+2] = (1.1*10^-10)/2 = 5.5*10^-11 M

then

BaSO4 = 5.5*10^-11 M